OpenStudy (anonymous):

State a rule that could be use to approximate the amout of time t needed to triple the amount of money in a savings account paying r percent interest compounded continuously.

OpenStudy (amistre64):

this one?

OpenStudy (amistre64):

A*e^(rt) = 3A e^(rt) = 3 t = ln(3)/r

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yes